Ok guys Im stuck again and need your help on this damn Cognitive Tudor program, the help and hints on it suck! Cant wait until I finish this class!
I need help to...Graph the equation x-3.1y=8.8 and 9x-9y=-37....I need to find their intersection again on the graph....Please someone help me again! Thanks!

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

This involves a lot of grunt work.
x-3.1y=8.8 and 9x-9y=-37
x=8.8+3.1y
9x-9y=-37
Plug that in for x and have fun.
x=-10.2593, y=-6.14815

Answer by Alan3354(69443)

dl the FREE graph software at
http://www.padowan.dk.com